OpenStudy (anonymous):

cos(x/5)sin(x/5)=1/2[sin(2x/5)] A. True B. False

OpenStudy (oblivious):

True By Trig identities cos(u)sin(v) = 1/2[sin(u+v)+sin(u-v)] =1/2[sin(x/5+x/5)+sin(x/5-x/5)] and since sin(0) is 0 this simplifies to 1/2[sin(2x/5)]